OpenStudy (anonymous):

Solve the equation. Check for extraneous solutions. log base of 4 (x+6) + log base of 4(x)= 2 do the logs base of 4 cancel out??

OpenStudy (anonymous):

log_4 (x + 6) + log_4 (x) = 2 Log properties. log_4 [x(x + 6)] = 2 4 to the power of both sides x² + 6x = 4^2 x² + 6x - 16 = 0 (x + 8)(x - 2) = 0 x = -8, 2 No real solutions to logarithms of neg numbers, so x = 2.